### v1

# Attention Retrieves, MLP Memorizes: Disentangling Trainable Components in the Transformer

Abstract
The Transformer architecture is central to the success of modern Large Language Models (LLMs), in part due to its surprising ability to perform a wide range of algorithmic tasks – including mathematical reasoning, memorization, and retrieval – using only gradient-based training on next-token prediction. While the core component of a Transformer is the self-attention mechanism, we question how much, and which aspects, of the performance gains can be attributed to it. To this end, we compare standard Transformers to variants in which either the multi-layer perceptron (MLP) layers or the attention projectors (queries and keys) are frozen at initialization. To further isolate the contribution of attention, we introduce MixiT – the Mixing Transformer – a simplified, principled model in which the attention coefficients are entirely random and fixed at initialization, eliminating any input-dependent computation or learning in attention. Surprisingly, we find that MixiT matches the performance of fully trained Transformers on various algorithmic tasks, especially those involving basic arithmetic or focusing heavily on memorization. For retrieval-based tasks, we observe that having input-dependent attention coefficients is consistently beneficial, while MixiT underperforms. We attribute this failure to its inability to form specialized circuits such as induction heads – a specific circuit known to be crucial for learning and exploiting repeating patterns in input sequences. Even more interestingly, we find that attention with frozen key and query projectors is not only able to form induction heads, but can also perform competitively on language modeling. Our results underscore the importance of architectural heterogeneity, where distinct components contribute complementary inductive biases crucial for solving different classes of tasks. [^1] Our code is publicly available at https://github.com/princeton-pli/MixiT.
1 Introduction
Transformers [48] have rapidly become the workhorse architecture in modern machine-learning systems, powering state-of-the-art models in language, vision, and scientific domains [15, 46, 21]. Their success is typically attributed to the self-attention mechanism, which allows every token to aggregate information from the entire sequence and has been linked to emergent abilities such as long-range retrieval, algorithmic reasoning, and in-context learning. Yet we lack a precise answer to a fundamental question: which degrees of freedom inside the Transformer are truly necessary for these behaviours, and which can be simplified away without harming performance?
Prior work has probed the internals of trained Transformers. Studies of attention maps consistently report the emergence of induction heads that copy information forward and enable in-context retrieval [36]. Another line of work has focused on whether and how Transformers’ capabilities can emerge [1, 40, 23] by studying synthetic datasets with Zhong and Andreas [57] showing that models with frozen self-attention but trainable embeddings still solve many algorithmic tasks. These results hint that different parts of the architecture are responsible for modeling different tasks.
A standard Transformer block, however, contains several interacting components: an attention block, with learnable and input-dependent queries, keys form the attention weights, and MLP blocks composed of fully connected layers. To exemplify the complexity of the interaction, notice that the attention weights can change their value both through changes in the queries/keys parameters, as well as as through the residual-stream representations that feed those projections. Hence, because all the components are trained together, it is hard to discern the contribution of each individual one to the solution of a given task. In this work, we study the role of each Transformer’s component by freezing different parts of the architecture to their value at initialization. In partiucular, we consider the following simplified variants:
  • Frozen-QK, preserves the conventional attention structure but freezes the query and key weight matrices, allowing only the value weight matrix to be learned.
  • Frozen-MLP, where the weight matrices of the MLP block are frozen to their initial value.
  • MixiT (Mixing Transformer), a Transformer variant in which the attention sub-layer is replaced by a fixed, randomly initialised mixing matrix. After initialisation the mixing weights are frozen, preventing the model from adapting interactions during training or conditioning them on input content. Learning therefore takes place exclusively in the embedding and MLP blocks and in residual branches.
Figure 1: Variants of the Llama Transformer model that we study.
Comparing these three simplified models with a fully-trainable Transformer makes it possible to separate the role of (i) learned token-dependent weights and (ii) the mere presence of a mixing operation, as well as (iii) understand the role of MLPs.
We compare across several categories of tasks: mathematical reasoning, sentiment classification, memorization, information retrieval and language modeling. This way, we can inspect whether and how the different parts of the architecture are able to solve primitive tasks that relate to a range of basic reasoning and memorization skills. Here we summarize our main contributions:
  • We find that, in all but information retrieval and language modeling, MixiT achieves performance comparable to the fully trained transformer and Frozen-QK. These results suggest that for a wide range of tasks, "attention mixing is all you need" regardless of its specific form, in the sense that learned query-key interactions are not required.
  • When tested on an information retrieval task, MixiT fails to achieve good performance, while both the fully trained transformer and Frozen-QK can solve it to close-to perfect accuracy. We show that while the alternatives do form induction heads, MixiT cannot do so by design, and this strongly correlates with the observed (in)ability to solve the task.
  • On pure memorization tasks, Frozen-MLP loses two-thirds of the storage capacity of the fully-trainable Transformer, while Frozen-QK and MixiT only a third. This suggests that the MLP layers are largely responsible for memorization, and input-dependent attention has a very little added contribution.
  • We present the surprising finding that Frozen-QK can perform competitively with the standard transformer on language modeling tasks. Indeed, Frozen-QK develops the ability to form specialized circuits such as induction heads during training.
Our results indicate that the Transformer architecture has some built-in inductive bias towards algorithmic abilities, namely the ability to form circuits. Furthermore, while for a variety of tasks no particular structure in the attention maps is needed, some learnable attention mechanism is required when flexibly copying tokens. As a consequence, due to the observed gap in language modeling, we speculate that the ability to solve associative recall and complex retrieval is a key ingredient to modern language models. More broadly, our results underpin the importance of having heterogeneous components that are more specialized in efficiently solving modular subtasks, enabling the model to express complex, wide-ranging behaviors.
2 Models Specification
We consider a decoder-only Transformer based on the widely used Llama architecture [47]. Given an input sequence $x\in\mathbb{R}^{V\times m}$, where $V$ is the vocabulary size and $m$ is the sequence length, embedded with a linear map to the hidden states $h_{0}=W_{emb}\,x$ where $W_{emb}\in\mathbb{R}^{V\times n}$, where $n$ is the width of the model. At its core, the transformer uses $L$ stacked modules alternating the causal multi-head self-attention layers and MLP layers. Each self attention head is defined as:
$\text{Attn}(h_{\ell})=W^{v}_{\ell}h_{\ell}\,\text{Softmax}\left(\frac{1}{\sqrt{n_{h}}}Q_{\ell}^{T}K_{\ell}\right)\,,\quad Q_{\ell}=W_{\ell}^{Q}h_{\ell}\,,\quad K_{\ell}=W_{\ell}^{K}h_{\ell}\,,$ (2.1)
where $W^{Q}_{\ell},W^{K}_{\ell},W^{v}_{\ell}\in\mathbb{R}^{n_{h}\times n}$ are the queries’ and keys’ weights, and the outputs across multiple heads are concatenated. The gated MLP layer is defined as:
$\text{MLP}(h_{\ell})=W^{D}_{\ell}\left(\phi(W_{\ell}^{G}h_{\ell})\odot W_{\ell}^{U}h_{\ell}\right)\,.$ (2.2)
where $W_{\ell}^{G},W_{\ell}^{U}\in\mathbb{R}^{{n_{m}\times n}}$ and $W^{D}_{\ell}\in\mathbb{R}^{{n\times n_{m}}}$. $n_{h}$ and $n_{m}$ are the dimensions of the queries/keys for each head and MLP projections. We also use skip connections in both blocks with a pre-normalization scheme [50] and causal masking. We apply rotary embeddings [43] to the queries and keys of each layer. As in common practice, we use $n_{m}=4n$ and $Hn_{h}=n$.
Models with Frozen Weights.
In the Frozen-QK model, we set $W^{Q}_{\ell},W^{K}_{\ell}$ to their value at initialization, and for the Frozen-MLP models we freeze $W^{D}_{\ell},W^{G}_{\ell},W^{U}_{\ell}$ for all layers.
MixiT – Random Static Attention.
We also design a model where the attention map itself is frozen and, to achieve that, input-independent. In the simplest case, this can be obtained by having a random matrix $M_{\ell}\in\mathbb{R}^{m\times m}$ entries with $\mathcal{N}(0,1/m)$ entries, where the factor of $1/m$ acts as a variance-preserving normalizer. To ensure a stable forward pass in terms of depth and width scaling, we follow the principles of attention shaping [34] and propose the following:
$\text{Attn}(h_{\ell})=W^{v}_{\ell}h_{\ell}\left(I+W^{M}_{\ell}-\bar{W}^{M}_{\ell}\right)\,,\quad W^{v}_{\ell,ij}\overset{iid}{\sim}\mathcal{N}(0,\tfrac{1}{n})\,,W^{M}_{\ell,ij}\overset{iid}{\sim}\mathcal{N}(0,\tfrac{1}{\sqrt{nm}})\,,$ (2.3)
where $W^{M}_{\ell}$ is frozen at initialization and $\bar{W}^{M}_{\ell}$ contains the column-wise empirical average of $W^{M}_{\ell}$, to ensure that each row sums up to 1. In the Appendix, , we show that this attention has a stable forward pass, in the sense that the kernel of the activations has a well-defined depth-and-width limit, converging to a stochastic differential equation (SDE) [28, 34]. When we adopt this architecture, all the weights, excluding the random attention matrix, are trainable. Notably, the following convergence result implies the stability of the forward pass, in particular ruling out the numerical degeneracy such as rank collapse and vanishing gradients [14, 33].
Theorem 2.1 (MixiT Covariance SDE).
Consider the MixiT recursion $h_{\ell+1}=\text{Attn}(h_{\ell})$ defined by Equation 2.3 at initialization. Then as the width $n$ and depth $d$ go to infinity with $\frac{d}{n}\to\bar{\tau}>0$, the upper triangular entries of the covariance matrix $\Phi_{\ell}=\frac{1}{n}h_{\ell}^{\top}h_{\ell}$ flattened to a vector in $\mathbb{R}^{m(m+1)/2}$ converges to the solution of the following SDE
$d\Phi_{\tau}=\left[\frac{1}{m}\operatorname{Tr}(\Phi_{\tau})-M(\Phi_{\tau})\right]\,d\tau+[\Sigma^{v}(\Phi_{\tau})+\Sigma^{M}(\Phi_{\tau})]^{1/2}\,dB_{\tau}\,,$ (2.4)
where $M(\Phi)=\frac{1}{m^{2}}\sum_{\alpha\beta=1}^{m}\Phi^{\alpha\beta}$ is the average over all entries, $B_{\tau}$ is a standard Brownian motion in $\mathbb{R}^{m(m+1)/2}$, $\Sigma^{v}(\Phi)^{\alpha\beta,\gamma\delta}=\Phi^{\alpha\gamma}\Phi^{\beta\delta}+\Phi^{\alpha\delta}\Phi^{\beta\gamma}$ and
$\Sigma^{M}(\Phi)^{\alpha\beta,\gamma\delta}=\delta_{\alpha\gamma}C(\Phi^{\bullet\beta},\Phi^{\bullet\delta})+\delta_{\alpha\delta}C(\Phi^{\bullet\beta},\Phi^{\bullet\gamma})+\delta_{\gamma\beta}C(\Phi^{\bullet\delta},\Phi^{\bullet\alpha})+\delta_{\beta\delta}C(\Phi^{\bullet\alpha},\Phi^{\bullet\gamma})\,,$ (2.5)
where $\delta_{\alpha\gamma}$ is the Kronecker delta, $C(\Phi^{\bullet\beta},\Phi^{\bullet\delta})=\frac{1}{m}\langle\Phi^{\bullet\beta},\Phi^{\bullet\delta}\rangle-\overline{\Phi}^{\bullet\beta}\overline{\Phi}^{\bullet\delta}$, $\Phi^{\bullet\beta}=[\Phi^{\alpha\beta}]_{\alpha=1}^{m}$ is the $\beta$-th column vector, and $\overline{\Phi}^{\bullet\beta}=\frac{1}{m}\sum_{\alpha}\Phi^{\alpha\beta}$ is the average.
The full proof can be found in the Appendix.
Positional embedding.
As the random attention matrix $I+\frac{1}{\sqrt{mn}}W^{M}_{\ell}-\bar{W}^{M}_{\ell}$ in MixiT does not depend on the input, the rotary positional embedding [43] standard in Llama models cannot be applied to MixiT, as rotary embeddings are added to learned key and query embeddings, which are not present in MixiT. Hence, we implement a learnable positional embedding for each token position in the sequence, and add it to the corresponding token embedding in the first layer.
3 Experiments
3.1 Tasks
We benchmark on a variety of tasks, covering categories including basic mathematical reasoning, memorization, sentiment classification, and language modeling. The basic mathematical reasoning and memorization tasks are based on tasks used in [57], with increased difficulty on some tasks to better reflect differences between architectures.
Decimal Addition.
For the decimal addition task, the model learns to add two integers with the same number of digits. We randomly sample 50,000 pairs of ten-digit numbers, and train the model to predict their sum. The test set consists of 4,000 such sequences not in the training set. An example is $1234567890+2345678901\to 3580246791$.
Needle in a Haystack (Retrieval).
Each training instance encodes a small randomly generated sequence of pairs followed by a single query, and the model is required to emit the value associated with that query one step later. We sample a sequence length $m\sim\mathcal{U}\{1,\dots,m_{\max}\}$, where $m_{\max}$ is the maximum sequence length. We then sample $m$ keys $\{k_{i}\}_{i=1}^{m}$ iid from the set $\{\frac{V}{2},\dots,V\}$ and $m$ values $\{v_{i}\}_{i=1}^{m}$ from $\{1,\dots,\tfrac{\text{V}}{2}-1\}$, where $V=256$ is the vocabulary size. The resulting keys are interleaved with their values to form the prefix $[(k_{1},\,v_{1}),\,(k_{2},\,v_{2}),\,\dots,\,(k_{m},\,v_{m})]$. A query key $k_{q}$ with $q\in[m]$ is chosen uniformly at random from the keys, and appended to the sequence. The goal is to predict the value corresponding to the query token. This task isolates retrieval ability, and probes associative recall. We sample 40,000 sequences for training, and 4,000 sequences for testing. Transformer-based models typically solve this task by forming induction heads [36].
Modular Addition.
This task evaluates the model’s ability to perform addition, modulo a given prime $p$. In our case we sample 40,000 pairs $(a,b)$ of integers, each within the range $[1,p]$, for $p=599$, and train the model to predict $a+b\text{ mod }p$.
Parentheses Balancing (Dyck-1).
[…]
Memorization.
[…]
Language Modeling.
To test the model’s ability to model natural language, we train the model to perform next-token-prediction. We use two datasets, Wikitext-103 [31] and Fineweb-edu [37]. Wikitext-103 consists of 1,801,350 cleaned-up Wikipedia articles, with test set of size 4358. And Fineweb-edu consists of top-quality entries collected from web-crawled data, focusing on educational text. We randomly sample 1,048,576 entries for training, and 4096 to test.
Sentiment Classification.
We use the Yelp polarity reviews dataset [56] to test each model variation’s ability to predict a review’s sentiment, i.e. whether a review is positive or negative.
3.2 Model training
For each task, we perform a grid search over a range of hyperparameters to train all model variations. The optimal hyperparameters are determined using a grid search specific for each task, which are detailed in the Appendix. All model variants are trained on one to four H100 GPUs, depending on task complexity.
4 Results
Figure 2: Retrieval accuracy as a function of the number of pairs in the sequence $m_{\max}$, which encodes task complexity. Having trainable queries and keys is necessary to solve the task (Standard and Frozen-MLP). Frozen-QK cannot solve the task for large $m_{\max}$, while MixiT has deteriorating performances.
Random static attention can perform certain algorithmic tasks.
Table 1 contains the main results on algorithmic and sentiment classification tasks with a large memorization focus.
[…]
Standard 98.58% 95.80% 100% 100% 90.55%
Frozen-QK 100% 97.38% 100% 100% 90.86%
MixiT 100% 96.17% 100% 100% 92.56%
Table 1: MixiT performance on algorithmic and sentiment classification tasks, many with a large memorization focus. As shown, both Frozen-QK and MixiT are able solve such tasks. They are competitive with, and can even be superior than, the standard transformer.
Trainable attention weights are necessary for retrieval.
We perform an experiment using the needle in a haystack dataset where the degree of complexity is controlled by the maximal number of pairs $m_{\max}$. We fix the embedding dimension $n=1024$ divided across $4$ heads with $n_{h}=256$, and fix 2 layers, which is sufficient to solve this task by using induction heads. We search for the optimal learning rate in the range $[5\cdot 10^{-5},5\cdot 10^{-3}]$. The results are shown in Figure 2. Notice how the while models with trainable queries and keys can solve the task across all the levels of difficulty, Frozen-QK has a drop in performances at $m_{\max}=40$ and MixiT struggles significantly after at $m_{\max}\geq 20$. These results provide compelling evidence that the MLP and embedding layers alone are insufficient when it comes to complex retrieval tasks.
Input-dependent attention patterns are required for language modeling.
Surprisingly, Frozen-QK comes close to the standard transformer in terms of perplexity, as shown in Table 2. This indicates that trainable attention weights are not always required for successful language modeling. Indeed, as Figure 3 shows, specialized circuits such as induction heads can form even in Frozen-QK. Less surprisingly, MixiT lags behind the standard Transformer, supporting the hypothesis that input-dependent learned attention patterns, such as induction heads, are necessary for language modeling, corroborating earlier works [36, 12].
Model \ Task Wikitext $\downarrow$ Fineweb-edu $\downarrow$
Standard 2.78 3.05
Frozen-QK 3.07 3.16
MixiT 3.73 4.08
Table 2: Performance on language modeling tasks, in terms of log perplexity. Frozen-QK comes surprisingly close in performance to the standard Transformer, despite having random static query and attention weights. MixiT lags behind the standard Transformer, supporting the hypothesis that input-dependent learned attention patterns, such as induction heads, are necessary for good language modeling.
Memorization.
We train an $L=2$, $n_{h}=4$ heads model on the memorization tasks for $10,000$ steps with a learning rate of $0.005$, for a consistent comparison across models. In Table 3 we report the accuracy and the storage capacity bits per parameters. We find that the standard Transformer stores $2.98$ bits per parameters, which is slightly higher than in previous works [1, 57]. Most of the drop occurs in the Frozen-QK model, with $1.13$ bits per parameters, while Frozen-MLP and MixiT have similar storage capabilities at $2.25$ and $2.18$, respectively. Note that these results are yielded in a setting where the accuracies are not saturated at 100%, to give an accurate representation of bits per parameter, rather than a lower bound. Hence they do not contradict the results in Table 1.
Model Memorization Accuracy $\uparrow$ Bits Per Parameter $\uparrow$ Trainable Parameters
Standard 100% 2.98 790400
Frozen-MLP 19% 1.13 394880
Frozen-QK 69% 2.25 724352
MixiT 67% 2.18 724736
Table 3: Standard Transformers outperform all the alternatives in terms of memorization capability, which suggests that MLPs and attention collaborate to remember knowledge. This provides further evidence for recent findings such as knowledge circuits [52] and query localization [10]. Freezing the MLPs causes the most performance drop, indicating that they are the biggest factor when it comes to memorization. Notice that MixiT has slightly more parameters than Frozen-QK because of additional trainable positional embeddings.
These results suggest that (1) the MLPs are largely responsible for memorization, however (2) there is a non-negligible additional contribution given by the integration of MLPs with learnable attention weights. This non-negligible additional contribution provides further evidence for recent findings such as knowledge circuits [52] and query localization [10], in that MLPs and attention collaborate to remember knowledge. In particular, the disproportionately large increase in the learned bits per parameter from Frozen-QK to the fully trained transformer, from 2.25 to 2.98, suggests that the gain in accuracy is more than what can be accounted for by a mere increase in learnable parameters.
Performance with respect to number of heads for MixiT.
For some tasks, we observe that increasing the number of attention heads in MixiT can notably improve performance, as demonstrated by the decimal addition and Dyck-1 parentheses balancing tasks in Table 4. Intuitively, since each attention head uses a different random attention matrix, more attention heads gives the learnable MLP components more diverse attention patterns to choose from based on the input, hence lessening the disadvantage of static attention. Note that purely increasing the number of heads, without increasing the hidden dimension, reaches diminishing returns, as the per-head embedding dimension decreases proportionally, restricting expressiveness.
Task \ Number of heads 4 16 64 256
Decimal addition 34.71% 34.70% 50.72% 91.87%
Dyck-1 77.93% 81.68% 89.38% 91.83%
Yelp sentiment classification 92.52% 92.56% 92.48% 91.72%
Table 4: Accuracy with respect to the number of heads on various tasks for MixiT. The hidden dimension is 512 for decimal addition and Dyck-1, and 1024 for Yelp sentiment classification. Increasing the number of heads increases the number of random attention matrices, giving learnable MLPs more diverse token mixing patterns to choose from based on the input, which can mitigate the disadvantage of static attention. However, a positive performance correlation does not appear in all tasks, such as in Yelp.
However, we do not observe this performance boost consistently across tasks. For instance, for Yelp sentiment classification, the performance is invariant with respect to the number of heads. This phenomenon remains interesting work for future study.
5 Discussion
Circuit learning and task separation.
As the attention matrix in MixiT is static and input-independent, MixiT cannot adapt to each input and form specific circuits such as induction heads [36]. Induction head circuits look for recent occurrences of the current token, and attends to the tokens that follow with increased probability. This allows the standard transformer to easily adapt to the input context in language modeling. Hence, it is no surprise that MixiT lags behind standard transformer for language modeling. It is perhaps more surprising that the perplexity comes close to that of the standard transformer.
The near-perfect performance on certain algorithmic tasks in Table 1 suggests that induction heads and other specialized input-dependent circuits are not required on these tasks, hence providing an axis for task separation.
Note that MixiT is able to perform well on the Yelp reviews dataset, despite the complexity of language used in reviews. This supports the hypothesis that sentiment can be largely judged by the collective token embeddings, as opposed to next-token prediction in language modeling tasks, which requires retrieving specific details from the context, a task induction heads are apt at.
Learnable attention is not required to form induction heads.
Interestingly, as shown in in Figure 3 and demonstrated by its performance, the Frozen-QK model can solve the retrieval task by forming induction heads.
Figure 3: The Frozen-QK model can solve the retrieval task by forming an induction head. In the first head, each token attends to the previous one; in particular, the query token 83 is attended by 256. In the head of the second layer, the correct token 256 is retrieved.
Role of MLPs in knowledge storage.
Previous works have highlighted the importance of MLPs in storing knowledge [13, 20, 19, 18, 53, 11, 30, 32]. They posit that specific facts from the training data are stored in specific knowledge neurons. These works support our findings, in that MLPs are crucial in knowledge memorization. However, our work does not prescribe knowledge localization, i.e. we don’t attribute memorizing specific facts to specific neurons.
Our work adds characterization on knowledge memorization in more recent works on knowledge circuits [52] and query localization [10], where MLPs and attention are found to collaborate on knowledge memorization, e.g. where attention selects the appropriate knowledge neurons depending on the query. Our work shows that even with static random attention weights, such as in Frozen-QK, attention and MLPs can still collaborate effectively, as evident in language modeling perplexities similar to that of the standard transformer, and the formation of specific input-dependent circuits. However, our results show that the role of MLPs in memorization outweighs that of attention, as evident by the fact that Frozen-MLP achieves much worse accuracy than Frozen-QK or MixiT (3).
Relation to random transformers.
Zhong and Andreas [57] studies the random transformer, wherein they train only the embedding and unembedding layers, and leave the intermediate layers fixed at random initialization. The random transformer was found to be able to perform nontrivial algorithmic tasks. There are several notable differences between MixiT and the random transformer. The principal difference is that the random attention matrix $I+W^{M}_{\ell}-\bar{W}^{M}_{\ell}$ in MixiT is input-independent, and hence cannot develop specialized circuits that adapt to the input, whereas such circuits are possible for the random transformer. Moreover, the MLPs in MixiT are trained, whereas they remain frozen in the random transformer. These differences allow us to study the roles different trained components play in the transformer architecture.
Furthermore, through in depth hyperparameter searches, we find that the random transformer does not scale well with respect to depth, confirming some of the original findings [57]. However, MixiT, with its carefully normalized random attention matrix designed to preserve signal propagation, does scale with respect to depth, suggesting that the random transformer suffers from signal propagation challenges and rank collapse without appropriate shaping [14, 33, 34].
Model \ Depth 2 8 16
MixiT 100% 100% 100%
Random transformer 100% 23.53% 22.88%
Table 5: Performance comparison between MixiT and the random transformer, with respect to number of layers, on the decimal addition task. The random transformer’s performance does not scale well with respect to depth, whereas proper attention matrix shaping helps MixiT scale with respect to depth.
Implications for architecture design.
Notably, some of these results strengthen the argument for empirical approaches for architecture design found in previous work [38, 8]. In particular, Poli et al. [38] uses performances in various synthetic tasks to design powerful hybrid architectures. Our results suggest that in certain circumstances some architectural components are entirely responsible for some basic operations (e.g. information retrieval) while in some other there is a more coordinated effort. Thus, overall this favours hybrid architectures with diverse basic components.
6 Related work
Static Attention.
Several studies explore simplified Transformers with frozen or random components. Notably, random Transformers with fixed layers but trainable embeddings can solve many algorithmic tasks [57]. Similarly, replacing attention with fixed matrices - as in Synthesizer [44] or FNet [27] — retains competitive performance on certain benchmarks, suggesting learned attention is not always necessary. However, retrieval tasks often require flexible, input-dependent attention to form induction-like circuits. While their attention is also random, unlike our MixiT model it is input-dependent and so does not isolate the specific tasks for which attention is and is not needed. We relate our results to their work in detail in Section 5. Other past work has studied properties of other randomly frozen or lightly trained models, e.g. convolutional networks [24, 41, 3], largely without focusing on specific tasks.
Stable Signal Propagation.
While there is a long line of work studying signal propagation in deep neural networks [42, 39, 26, 51], it was only more recently that Martens et al. [29] introduced the concept that modifying activation functions can significantly improve stability of signal propagation, and leading to rapid training of large scale vision models without using normalization and skip connections [55]. This was later understood to yield a stable scaling limit, which is characterized by an SDE of the covariance matrix [28]. The covariance SDE framework is then used to understand how to design and shape general non-linearities like the Transformer self-attention [34], resolving the rank collapse issue [14, 33] and has shown strong performance despite a much simplified Transformer block [22]. Our theoretical result Theorem 2.1 also follows from this framework.
Modular Tasks.
Several works have investigated the capabilities of Transformers in the classes of tasks analyzed here, including arithmetic [35]. The role of feedforward layers in memorization in the Transformer architecture has been studied in Geva et al. [17] and their inductive biases and scaling properties have been scrutinized [5]. In this context, our work shows their relevance in conjuction with trainable or fixed attention. Orthogonal to our work, memorization has also been studied to understand generalization in neural networks [54, 4, 2].
Mechanistic Interpretability.
Other closely related work is in the mechanistic interpretability, which aims to understand LLMs via examining and modifying their internals [6]. Closely related is work related to identifying and understanding the behavior of induction heads [36, 16, 7] and in-context learning [9]. Our work demonstrates that the performance separation between input-dependent and input-independent attention is largely driven by the latter’s inability to form induction heads. Finally, Meng et al. [30] which reverse-engineer how different Transformer components support behaviors like memorization, retrieval, and generalization.
Efficient Attention.
A last area that has seen significant effort at understanding attention is that of efficient Transformers [45]. While we study what happens when attention is replaced with a fixed input-independent matrix, this field has studied various useful aspects of the attention matrix such as attention sinks [49] and compression [25]. As our work demonstrates that for many tasks the full power of input-dependent attention is not needed, it may have its own implications for efficiency, e.g. by removing the need for the KV-cache.
7 Conclusion
In this work, we isolated the roles of the architecture’s two trainable components by selectively freezing parameters and introducing MixiT, whose attention maps are fixed and random. Fixed attention suffices for many algorithmic tasks. However, adaptive attention is essential for associative recall and language modelling, as MixiT fails on long-range retrieval and yields higher perplexity. However, memorization resides in the MLPs, with attention contributing significantly less. Together, these results support a heterogeneity hypothesis: MLPs and learnable attention mechanism are required for the full behavioural range of modern large language models.
Limitations and future work
Our work studied basic algorithmic, mathematical reasoning, and language modeling tasks. It is an exciting future direction to extend our study to more complex tasks such as human reasoning, which demands a wide-ranging skillset that does not fall neatly on one side of the retrieval-memorization spectrum. We used relatively small-scale models; extending the analysis to larger networks and datasets will test whether the observed separation of concerns persists at a larger scale. Future work might study hybrid training schedules in which only a subset of architectural modules remain trainable – or are gradually unfrozen – which may strike an even better accuracy–efficiency trade-off. Finally, probing how these architectural choices interact with emerging interpretability and safety techniques constitutes an exciting avenue for further research.
Appendix
Appendix A Proof of Theorem 2.1
Theorem (MixiT Covariance SDE).
Consider the MixiT recursion $h_{\ell+1}=\text{Attn}(h_{\ell})$ defined by Equation 2.3 at initialization. Then as the width $n$ and depth $d$ go to infinity with $\frac{d}{n}\to\bar{\tau}>0$, the upper triangular entries of the covariance matrix $\Phi_{\ell}=\frac{1}{n}h_{\ell}^{\top}h_{\ell}$ flattened to a vector in $\mathbb{R}^{m(m+1)/2}$ converges to the solution of the following SDE
$d\Phi_{\tau}=\left[\frac{1}{m}\operatorname{Tr}(\Phi_{\tau})-M(\Phi_{\tau})\right]\,d\tau+[\Sigma^{v}(\Phi_{\tau})+\Sigma^{M}(\Phi_{\tau})]^{1/2}\,dB_{\tau}\,,$ (A.1)
where $M(\Phi)=\frac{1}{m^{2}}\sum_{\alpha\beta=1}^{m}\Phi^{\alpha\beta}$ is the average over all entries, $B_{\tau}$ is a standard Brownian motion in $\mathbb{R}^{m(m+1)/2}$, $\Sigma^{v}(\Phi)^{\alpha\beta,\gamma\delta}=\Phi^{\alpha\gamma}\Phi^{\beta\delta}+\Phi^{\alpha\delta}\Phi^{\beta\gamma}$ and
$\Sigma^{M}(\Phi)^{\alpha\beta,\gamma\delta}=\delta_{\alpha\gamma}C(\Phi^{\bullet\beta},\Phi^{\bullet\delta})+\delta_{\alpha\delta}C(\Phi^{\bullet\beta},\Phi^{\bullet\gamma})+\delta_{\gamma\beta}C(\Phi^{\bullet\delta},\Phi^{\bullet\alpha})+\delta_{\beta\delta}C(\Phi^{\bullet\alpha},\Phi^{\bullet\gamma})\,,$ (A.2)
where $\delta_{\alpha\gamma}$ is the Kronecker delta, $C(\Phi^{\bullet\beta},\Phi^{\bullet\delta})=\frac{1}{m}\langle\Phi^{\bullet\beta},\Phi^{\bullet\delta}\rangle-\overline{\Phi}^{\bullet\beta}\overline{\Phi}^{\bullet\delta}$, $\Phi^{\bullet\beta}=[\Phi^{\alpha\beta}]_{\alpha=1}^{m}$ is the $\beta$-th column vector, and $\overline{\Phi}^{\bullet\beta}=\frac{1}{m}\sum_{\alpha}\Phi^{\alpha\beta}$ is the average.
Proof.
Firstly, we recall that based on Li et al. [28], the linear network covariance matrix $\Phi_{\ell}=\frac{1}{n}h_{\ell}^{\top}h_{\ell}$ for the recursion $h_{\ell+1}=W^{v}_{\ell}h_{\ell}$ for $W^{v}_{\ell,ij}\sim\mathcal{N}(0,\tfrac{1}{n})$ satisfies the Markov chain
$\Phi_{\ell+1}=\Phi_{\ell}+\frac{\Sigma^{v}(\Phi_{\ell})^{1/2}\xi_{\ell}}{\sqrt{n}}\,,$ (A.3)
where $\xi_{\ell}$ is a zero mean and identity covariance random variable, and the diffusion coefficient is $\Sigma^{v}(\Phi)^{\alpha\beta,\gamma\delta}=\Phi^{\alpha\gamma}\Phi^{\beta\delta}+\Phi^{\alpha\delta}\Phi^{\beta\gamma}$. Therefore, it is sufficient to isolate the contribution of the mixing component alone, and we will add the effect of the two components.
To this end, we consider the equivalent recursion
$h_{\ell+1}=h_{\ell}\left(I_{m}+\tfrac{1}{\sqrt{nm}}(W^{M}_{\ell}-\bar{W}^{M}_{\ell})\right)\,,$ (A.4)
where we consider $W^{M}_{\ell,ij}\sim\mathcal{N}(0,1)$ instead of $\mathcal{N}(0,\tfrac{1}{nm})$ due to the pre-factor, and $\bar{W}^{M}_{ij}=\tfrac{1}{m}\sum_{k=1}^{m}W^{M}_{kj}$ replaces each entry by its corresponding column average.
Next, we will observe that $\Phi_{\ell}$ satisfies a straight forward recursion
$\Phi_{\ell+1}$ $=\frac{1}{n}h_{\ell+1}^{\top}h_{\ell+1}$ (A.5)
$=\left(I_{m}+\tfrac{1}{\sqrt{nm}}(W^{M}_{\ell}-\bar{W}^{M}_{\ell})\right)^{\top}\Phi_{\ell}\left(I_{m}+\tfrac{1}{\sqrt{nm}}(W^{M}_{\ell}-\bar{W}^{M}_{\ell})\right)$
$=\Phi_{\ell}+\tfrac{1}{\sqrt{nm}}\left[\left(W^{M}_{\ell}-\bar{W}^{M}_{\ell}\right)^{\top}\Phi_{\ell}+\Phi_{\ell}\left(W^{M}_{\ell}-\bar{W}^{M}_{\ell}\right)\right]$
$\quad\quad+\tfrac{1}{nm}\left(W^{M}_{\ell}-\bar{W}^{M}_{\ell}\right)^{\top}\Phi_{\ell}\left(W^{M}_{\ell}-\bar{W}^{M}_{\ell}\right)\,,$
which naturally separates itself into the diffusion and drift components via the coefficient scale of $\tfrac{1}{\sqrt{nm}}$ and $\tfrac{1}{nm}$ respectively.
We will compute the drift term next. Here we will drop some super and subscripts to reduce clutter, and write
$\sum_{\alpha,\beta=1}^{m}\mathbb{E}_{\ell}(W-\bar{W})^{\alpha\gamma}\Phi^{\alpha\beta}(W-\bar{W})^{\beta\delta}$ $=\sum_{\alpha\beta}\Phi^{\alpha\beta}\mathbb{E}_{\ell}\left[W^{\alpha\gamma}W^{\beta\delta}-\frac{1}{m^{2}}\sum_{\mu\nu}W^{\mu\gamma}\bar{W}^{\nu\delta}\right]$ (A.6)
$=\sum_{\alpha\beta}\Phi^{\alpha\beta}(\delta_{\alpha\beta}\delta_{\gamma\delta}-\frac{1}{m^{2}}\sum_{\mu\nu}\delta_{\mu\nu}\delta_{\gamma\delta})$
$=\delta_{\gamma\delta}\sum_{\alpha\beta}\Phi^{\alpha\beta}\left(\delta_{\alpha\beta}-\frac{1}{m}\right)\,,$
where $\mathbb{E}_{\ell}[\,\cdot\,]=\mathbb{E}[\,\cdot\,|\mathcal{F}_{\ell}]$ and $\mathcal{F}_{\ell}=\sigma(\{h_{k}\}_{k\leq\ell})$, which translates to the final drift of
$\frac{1}{n}\left(\frac{1}{m}\operatorname{Tr}(\Phi)-M_{\Phi}\right)I_{n}\,,$ (A.7)
where $M_{\Phi}=\frac{1}{m^{2}}\sum_{\alpha\beta}\Phi^{\alpha\beta}$ is the average over all entries.
To calculate a single entry of the diffusion coefficient $\Sigma(\Phi)^{\alpha\beta,\gamma\delta}$, we will write $\widetilde{W}=W-\bar{W}$ and compute
$\Sigma^{M}(\Phi)^{\alpha\beta,\gamma\delta}$ (A.8)
$=\frac{1}{m}\sum_{\mu,\nu=1}^{m}\mathbb{E}_{\ell}(\widetilde{W}^{\mu\alpha}\Phi^{\mu\beta}+\Phi^{\alpha\mu}\widetilde{W}^{\mu\beta})(\widetilde{W}^{\nu\gamma}\Phi^{\nu\delta}+\Phi^{\gamma\nu}\widetilde{W}^{\nu\delta})$
$=\frac{1}{m}\sum_{\mu,\nu}\mathbb{E}_{\ell}\left[\widetilde{W}^{\mu\alpha}\Phi^{\mu\beta}\widetilde{W}^{\nu\gamma}\Phi^{\nu\delta}+\widetilde{W}^{\mu\alpha}\Phi^{\mu\beta}\Phi^{\gamma\nu}\widetilde{W}^{\nu\delta}+\Phi^{\alpha\mu}\widetilde{W}^{\mu\beta}\widetilde{W}^{\nu\gamma}\Phi^{\nu\delta}+\Phi^{\alpha\mu}\widetilde{W}^{\mu\beta}\Phi^{\gamma\nu}\widetilde{W}^{\nu\delta}\right]\,.$
At this point we focus on one term and compute
$\mathbb{E}_{\ell}\,\widetilde{W}^{\mu\alpha}\widetilde{W}^{\nu\beta}$ $=\mathbb{E}_{\ell}\,(W^{\mu\alpha}-\bar{W}^{\mu\alpha})(W^{\nu\beta}-\bar{W}^{\nu\beta})$ (A.9)
$=\mathbb{E}_{\ell}\,(W^{\mu\alpha}W^{\nu\beta}-W^{\mu\alpha}\bar{W}^{\nu\beta}-\bar{W}^{\mu\alpha}W^{\nu\beta}+\bar{W}^{\mu\alpha}\bar{W}^{\nu\beta})\,,$
$\delta_{\alpha\gamma}$ is the Kronecker delta, and we can separate further then compute
$\mathbb{E}_{\ell}\,W^{\mu\alpha}W^{\nu\beta}$ $=\delta_{\mu\nu}\delta_{\alpha\beta}\,,$ (A.10)
$\mathbb{E}_{\ell}\,W^{\mu\alpha}\frac{1}{m}\sum_{\nu^{\prime}=1}^{m}W^{\nu^{\prime}\beta}$ $=\frac{1}{m}\sum_{\nu^{\prime}}\delta_{\mu\nu^{\prime}}\delta_{\alpha\beta}=\frac{1}{m}\delta_{\alpha\beta}\,,$
$\mathbb{E}_{\ell}\,\bar{W}^{\mu\alpha}W^{\nu\beta}$ $=\frac{1}{m}\delta_{\alpha\beta}\,,$
$\mathbb{E}_{\ell}\,\bar{W}^{\mu\alpha}\bar{W}^{\nu\beta}$ $=\frac{1}{m^{2}}\sum_{\mu^{\prime},\nu^{\prime}=1}^{m}\delta_{\mu^{\prime}\nu^{\prime}}\delta_{\alpha\beta}=\frac{1}{m}\delta_{\alpha\beta}\,.$
This implies
$\mathbb{E}_{\ell}\,\widetilde{W}^{\mu\alpha}\widetilde{W}^{\nu\beta}=\delta_{\mu\nu}\delta_{\alpha\beta}-\frac{1}{m}\delta_{\alpha\beta}=(\delta_{\mu\nu}-\tfrac{1}{m})\delta_{\alpha\beta}\,.$ (A.11)
At this point, we return to calculating $\Sigma^{M}(\Phi)^{\alpha\beta,\gamma\delta}$ and write
$\Sigma^{M}(\Phi)^{\alpha\beta,\gamma\delta}$ $=\frac{1}{m}\sum_{\mu\nu}(\delta_{\mu\nu}-\tfrac{1}{m})\delta_{\alpha\gamma}\Phi^{\mu\beta}\Phi^{\nu\delta}+(\delta_{\mu\nu}-\tfrac{1}{m})\delta_{\alpha\delta}\Phi^{\beta\mu}\Phi^{\gamma\nu}$ (A.12)
$\quad\quad+(\delta_{\mu\nu}-\tfrac{1}{m})\delta_{\gamma\beta}\Phi^{\mu\delta}\Phi^{\alpha\nu}+(\delta_{\mu\nu}-\tfrac{1}{m})\delta_{\beta\delta}\Phi^{\alpha\mu}\Phi^{\gamma\nu}$
$=\delta_{\alpha\gamma}C(\Phi^{\bullet\beta},\Phi^{\bullet\delta})+\delta_{\alpha\delta}C(\Phi^{\bullet\beta},\Phi^{\bullet\gamma})+\delta_{\gamma\beta}C(\Phi^{\bullet\delta},\Phi^{\bullet\alpha})+\delta_{\beta\delta}C(\Phi^{\bullet\alpha},\Phi^{\bullet\gamma})\,,$
where $C(\Phi^{\bullet\beta},\Phi^{\bullet\delta})=\frac{1}{m}\langle\Phi^{\bullet\beta},\Phi^{\bullet\delta}\rangle-\overline{\Phi}^{\bullet\beta}\overline{\Phi}^{\bullet\delta}$, $\Phi^{\bullet\beta}=[\Phi^{\alpha\beta}]_{\alpha=1}^{m}$ is the $\beta$-th column vector, and $\overline{\Phi}^{\bullet\beta}=\frac{1}{m}\sum_{\alpha}\Phi^{\alpha\beta}$ is the average.
To complete the proof, we will invoke the Markov chain convergence to SDE results in the Skorohod topology, see for example Li et al. [28, Proposition A.6], which gives us the desired result.
Appendix B Additional Experimental Details
B.1 Hyperparameter selection
We select for the optimal hyperparameters for each task using a grid search. The search ranges for each task are as follows, determined a priori depending on task complexity, e.g. language modeling is inherently more complex than memorization.
Task \ Hyperparameter # Layers Hidden dimension # Heads Learning rate Batch size
Algorithmic [2, 4, 8] [512, 1024] [4, 8, 16, 64] [1e-3, 5e-4, 1e-4] [128, 256, 512]
Retrieval [2, 4, 8] [512, 1024] [4, 8, 16, 64] [1e-3, 5e-4, 1e-4] [256, 512, 1024]
Memorization [2, 4, 8] [512, 1024] [4, 8, 16, 64] [1e-3, 5e-4, 1e-4] [256, 512, 1024]
Yelp [4, 8] [512, 1024] [4, 8, 16, 64] [1e-3, 5e-4, 1e-4] [256, 512, 1024]
Language modeling [8, 12] [512, 1024] [4, 8, 16] [1e-3, 5e-4, 1e-4] [256, 512, 1024]
Table 6: Hyperparameter ranges used during grid search, for all architectures. Algorithmic tasks include decimal addition, Dyck-1 parentheses balancing, and modular addition.
We use a sequence length of 256 for Yelp sentiment classification and language modeling tasks. Language modeling tasks are trained for 40,000 steps.
Task \ Hyperparameter # Layers Hidden dimension # Heads Learning rate Batch size
Decimal Addition 8 512 64 1e-3 128
Dyck-1 4 512 64 1e-3 512
Modular addition 2 512 32 1e-3 256
Retrieval 2 1024 4 1e-4 1024
Memorization 2 1024 4 1e-3 256
Yelp 4 1024 16 5e-4 256
Wikitext 12 512 8 5e-4 512
Fineweb-edu 12 512 8 5e-4 512
Table 7: Optimal hyperparameters selected for MixiT.
Appendix C Further Discussion
This section contains further discussion on experiments and results.
Improved throughput for MixiT and Frozen-QK
Leads to a notable improvement in training time for language modeling. For instance, on the Fineweb-edu dataset, with the same architecture and hyperparameters on the same infrastructure, Frozen-QK trains 1267.1 samples per second on average, whereas the standard transformer trains 1022.8 samples per second on average, while achieving similar log perplexities, 3.05 for standard and 3.15 for Frozen-QK. This represents a 23.9% improvement in training throughput, leading to a 23.9% speedup in terms of wall clock time. MixiT trains even faster, with 1349.0 training samples per second, or a 32.0% improvement in throughput. However, MixiT comes with noticeable degradation in perplexity.